1. A personal director for corporation has hired eleven new engineers. If three different positions are open at Palawan plan, in how many ways can he fill the position?
2. A man is in the process of buying a new car. he has choices of four engine makes, 5 body styles and 12 colors. how many choices of diiferent cars does he have?

Answer by stanbon(75887)

1. A personal director for corporation has hired eleven new engineers. If three different positions are open at Palawan plan, in how many ways can he fill the position?
---
Ans: 11P3 = 11!/(11-3)! = 11*10*9 = 11*90 = 990 ways
-----------------------------
2. A man is in the process of buying a new car. he has choices of four engine makes, 5 body styles and 12 colors. how many choices of diiferent cars does he have?
------
Ans: General counting Rule:
# of cars = 4*5*12 = 4*60 = 240
-----
Cheers,
Stan H.
===================